A gas has a volume of 2.90 cm3 at 226 K. What would the volume be if the temperature were changed to 23.0 °C?
1
Expert's answer
2020-05-19T08:58:30-0400

P*V/T=P0*V0/T0

V=T*P0*V0/P*T0=296*2.90*P0/226*P=3.798*P0/P

if VT=const, V=3.8 mL